OpenStudy (anonymous):

Which is the graph functions f(x) = - 1/2(x + 2)^2 + 4?

OpenStudy (er.mohd.amir):

y2=4ax is parabola here your eq. same like it so graph is like parabola but vertex and focus is change find it to reduce to y^2 = 4 a x

OpenStudy (triciaal):

in vertex form y = -a(x-h)^2 + k parent function y = x^2 center (0, 0) (h,k) a = 1 here a = - reflection across x-axis a = 1/2 1/2 the y value center x + 2 moved 2 places left k + 4 = y moved 4 places up

OpenStudy (triciaal):